Part D: Problem 29. Consider a O.15 mol/L solution of NaOH a) Calculate the hydroxide ion concentretion. b) Calculate the pOH c) Calculate the pH d) Calculate the hydrogen ion concentration
Answer #1

29(a) In calculation of pH or pOH of acids and bases Normality should be considered, but in the question molarity is given. Let's convert molarity into normality.

Normality=Molarity × Valency

Here valency of NaOH is 1

Normality=0.15×1 = 0.15

This gives the concentration of [OH-]=0.15mol/l

(b) pOH = [-log(OH-)]

pOH = [-log(0.15)]

pOH = 0.82

(c) pH+pOH = 14

pH+0.823 = 14

pH = 13.18

(d) pH = [-log(H+)]

[H+] = 10-pH

[H+] = 10-13.18

[H+] = 6.607

Thank you.